OpenStudy (anonymous):

A bag contains 7 red marbles, 12 orange marbles, 18 yellow marbles, and 5 purple marbles. A marble is drawn at random from the bag. What is the probability that the marble drawn will not be orange?

OpenStudy (anonymous):

A.2/7 B.1/12 C.5/7 D.1/30

OpenStudy (y1066380):

Total number of marbles=7+12+18+5=42 Number of orange marbles=12 So, by using the equation "favorable outcome / possible outcome" You will have 12/42=2/7 You will get the ans A.

OpenStudy (anonymous):

Add the total amount of marbles

OpenStudy (anonymous):

Thanks to both of yous :)